What do you understand by temporal and spatial coherence of light? Is it true that observable interference fringe pattern cannot be obtained by using two incoherent sources in an interference experiment? Justify your answer.
1
Expert's answer
2020-02-11T10:14:02-0500

Temporal Coherence is a measure of the correlation of light wave’s phase at different points along the direction of propagation – it tells us how monochromatic a source is.


Spatial Coherenceis a measure of the correlation of a light wave’s phase at 

different points transverse to the direction of propagation - it tells us how uniform the phase of a wavefront is.


Only coherent waves can interfere. The distribution of light intensity in space is given by equation


"\langle E^2\rangle=\langle E^2_1\rangle+\langle E^2_2\rangle+2\langle \overrightarrow{E_1}\cdot \overrightarrow{E_2}\rangle" or "I=I_1+I_2+2\sqrt{I_1I_2}\cos\Delta\phi"


For incoherent waves


"\langle \overrightarrow{E_1}\cdot \overrightarrow{E_2}\rangle=0"


And "\langle E^2\rangle=\langle E^2_1\rangle+\langle E^2_2\rangle" or "I=I_1+I_2".